The length of a rectangle is twice the width. If the length is increased by 4 inches and the width is diminished by 1 inch, a new rectangle is formed whose perimeter is 198 inches. Fine the dimensions of the original rectangle
Answer by lwsshak3(6469)

The length of a rectangle is twice the width. If the length is increased by 4 inches and the width is diminished by 1 inch, a new rectangle is formed whose perimeter is 198 inches. Fine the dimensions of the original rectangle
..
let x=original length=2y
let y=original width
length increased by 4=x+4
width deminished by 1=y-1
Perimeter=2*length+2*width=198
2(x+4)+2(y-1)=198
2x+8+2y-2=198
substitute x for 2y
2x+8+x-2=198
3x=192
x=64 inches
y=x/2=64/2=32 inches
Ans:
Dimensions of original triangle: 64 by 32 inches